Write an equation of the line containing the given point and parallel to the given line. Express your answer in the for y=mx+b. (6,7); x+5y=4
Answer by mananth(16949)

(6,7);
x+5y=4
5y=-x-4
y=(-x/5)-(4/5)
..
compare with y= mx+b
m=(-1/5)
the slope of the required will be the same since line is parallel
Plug (6,7) & (-1/5) to get b
in y= mx+b
7=(-1/5)*6+b
7+6/5=b
b=41/5
The equation of line
y=(-1/5)x-41/5

...